# TTQA-RS- A break-down prompting approach for Multi-hop Table-Text Question Answering with Reasoning and Summarization

### v1

@@ Abstract
Question answering (QA) over tables and text has gained much popularity over the years. Multi-hop table-text QA requires multiple hops between the table and text, making it a challenging QA task. Although several works have attempted to solve the table-text QA task, most involve training the models and requiring labeled data. In this paper, we have proposed a model - “TTQA-RS: A break-down prompting approach for Multi-hop Table-Text Question Answering with Reasoning and Summarization” [^1] Link to code in Github: https://github.com/jayetri/TTQA-RS. Our model uses augmented knowledge including table-text summary with decomposed sub-question with answer for a reasoning-based table-text QA. Using open-source language models our model outperformed all existing prompting methods for table-text QA tasks on existing table-text QA datasets like HybridQA and OTT-QA’s development set. Our results are comparable with the training-based state-of-the-art models, demonstrating the potential of prompt-based approaches using open-source LLMs. Additionally, by using GPT-4 with LLaMA3-70B, our model achieved state-of-the-art performance for prompting-based methods on multi-hop table-text QA.
@@ 1 Introduction
Over the years, several works have attempted to solve this task. But the majority of these works have used supervised-training, requiring a large amount of labeled data (Chen et al., 2020b; Sun et al., 2021; Wang et al., 2022; Eisenschlos et al., 2021; Feng et al., 2022; Kumar et al., 2023; Chen et al., 2020a; Li et al., 2021). In this paper, we have proposed a prompting-based approach while using open-source large language models (LLMs) for multi-hop table-text QA.
@@
In this paper, we introduced a framework - TTQA-RS, a reasoning-based prompting approach for table-text QA that despite CoT’s shortcomings on small-parameter models, we were able to reduce the hallucinations on open-source small models (i.e. we obtained a 6% increase in exact match score compared to the baseline CoT model for the HybridQA’s test set). Furthermore, our proposed model was able to beat the state-of-the-art model - S3HQA’s CoT prompting with GPT 3.5 results (Lei et al., 2023) on HybridQA dataset. By beating their model’s performance, we have shown the potential for smaller LLMs in multi-hop table-text QA.
For our experiments, we have used HybridQA dataset and OTT-QA’s development set. OTT-QA is an extension of the HybridQA dataset. Similar to the HybridQA dataset, the OTT-QA dataset is also constructed using questions based on Wikipedia tables and text. But unlike the HybridQA dataset, the test set of the OTT-QA dataset does not have hyperlinks in the table cells that can be linked to the Wikipedia text. Hence, the OTT-QA’s test set is more challenging. Existing models including ours use a retriever-reader framework for table-text QA. In this paper, we narrowed our focus to the reader of the table-text QA task. Our goal is to develop a prompting strategy for the table-text QA reader that can work even with smaller LLMs. The task of linking tables and text passages for open-domain QA is out of scope of this paper. The development set of OTT-QA, similar to the HybridQA dataset, already has hyperlinks in the table cells linking to the wiki text. In the future, we plan to extend our approach to linking the table and text for cases when hyperlinks are absent in the table cells.
The TTQA-RS model breaks down the table-text QA problem into multiple steps. In the HybridQA and the OTT-QA dataset, the questions require multiple steps of reasoning over table and text to answer. The TTQA-RS model generates the sub-questions that can help in answering the complex questions. It also generates the summary of the table and text, which is in turn used for the table-text QA of the original questions. Breaking down the complex multi-hop QA problem into simple, smaller steps can help boost the model’s overall performance. Furthermore, LLMs struggle with multi-level reasoning in a single step. So, breaking down the multi-hop QA problem along with providing an augmented information including the table-text summary can improve the performance of multi-hop table-text QA tasks using small open-source LLMs. In Figure 1, we show an example of a question from multi-hop QA that uses standard prompting, CoT, and the TTQA-RS approach for multi-hop QA.
@@ 2 Related Works
Multi-hop table-text QA can be a complex task as it requires multiple hops between the table and text to answer the questions. S3HQA (Lei et al., 2023) and MFORT-QA (Guan et al., 2024) are the only two existing models as per our knowledge that use in-context learning for multi-hop table-text QA. The S3HQA model has demonstrated table-text QA task using the Hybrid-QA dataset, whereas MFORT-QA has used the OTT-QA dataset. The S3HQA model uses a three-step method - a retriever with refinement training, a hybrid selector, and a generation-based reasoner with GPT 3.5 for the hybrid table-text QA task. MFORT-QA uses the Chain-of-thought (CoT) method to break down complex questions into smaller sub-questions, and uses Retrieval Augmented Generation to extract more context. Similar to the MFORT-QA model, we also break down complex questions into smaller sub-questions. With the complexity of the multi-hop QA task broken down into smaller questions, LLMs are in turn working on a smaller problem and perform better as single-step reasoners. Our model - TTQA-RS, additionally generates a summary using the retrieved table rows and passages. Then, for table-text question answering (QA), it uses the generated summary, the predicted entity type of the answer, and the generated sub-questions along with the answer.
Figure 2: An overview of TTQA-RS framework. The dashed lines represent the reader for the table-text QA model.
@@ 3.1 System Overview
The TTQA-RS model uses a retriever-reader model. Our reader breaks down the table-text QA problem into five steps - (1) Summary generation using retrieved tables rows and passages, (2) Question decomposition, (3) Entity type prediction of the expected answer, (4) Table-text QA of independent sub-question, and (5) Table-text QA of the original question. Figure 2 shows an overview of the TTRS-QA framework. The following subsections describe the TTQA-RS framework’s retriever and reader in detail.
@@ 3.2 Retriever
The function of the retriever is to extract relevant rows and passages from the text linked to the table cells using hyperlinks. For the HybridQA dataset, we have used S3HQA model’s (Lei et al., 2023) table retriever to extract the relevant row(s) from the table, and HYBRIDER’s (Chen et al., 2020b) text-retriever to extract the relevant information from the linked passages. S3HQA’s row retriever uses refinement training to train the retriever model. The tables contain hyperlinks to Wikipedia text. So, the passages linked to the retrieved rows are collected to form a pool. The passage retriever contains an ensemble retriever of TF-IDF retriever with longest-substring retriever and selects passages with cosine distance less than a certain threshold.
For experiments on OTT-QA’s development set, we don’t use any table retriever, i.e. we only use HYBRIDER’s text retriever. The text linked to the table rows is extracted and then filtered using HYBRIDER’s text retriever.
@@ 3.3.1 Table-text Summarization
This is the first step of the reader model. The retrieved table rows are flattened with a delimiter separating the rows and columns of the table. The retrieved rows and passages are used to generate summaries of the table and text. We used zero-shot learning with LLaMA 3-70B model to generate the summaries. In Appendix B we have shown an example of a table-text summarization prompt.
Figure 3: Example of our approach using TTQA-RS model
@@ 3.3.2 Question decomposition
In the next step, we break down the questions and identify the sub-questions, such that the answer of one sub-question can aid in answering the original complex question. From here onwards, we will refer to the sub-question that can be answered first as the “independent sub-question”. Let’s take the example in Figure 3. The complex question - "What was the release date of the game which Andrew Voss provided commentary on ?" can be broken down into sub-questions. The independent sub-question for this question is - "Which game has Andrew Voss provided commentary on?". The answer to this sub-question is "Rugby League 3". This can be used to simplify the original complex question to the following - " What was the release date of Rugby League 3?". Thus, including the information about the independent sub-question and the sub-answer helps to reduce the complexity of the multi-hop task. Identifying the independent sub-question and breaking down the complex multi-hop QA problem helps to reduce the complexity of the problem, and in turn, boosts the accuracy of the model. We use in-context learning with LLaMA3-70B model to generate the independent sub-questions for the given complex queries.
@@ 4.1 Datasets
OTT-QA (Chen et al., 2020a) is an open-domain multi-hop table-text QA dataset. For our experiments, we only use the development set of the OTT-QA dataset which contains the hyperlinks linking the table and the text (unlike OTTQA’s test set).
@@ 4.3 Baseline Models
Standard prompting - For the baseline standard prompting model, we used the same retriever as in TTQA-RS model, (i.e. HYBRIDER’s (Chen et al., 2020b) passage retriever with S3HQA’s (Lei et al., 2023) table retriever for the HybridQA dataset). For experiments on OTT-QA’s dev set, we don’t use any table-retriever, i.e. we only use the HYBRIDER’s passage retriever. For the reader, we performed in-context learning with standard prompting (Brown et al., 2020) for the QA task.
Chain of Thought Prompting (CoT) - Similar to the standard prompting baseline model, the CoT baseline model uses the same retriever as the TTQA-RS model. The reader uses in-context learning with CoT prompting Wei et al. (2022).
@@ 5.1 Main Results
Train MuGER2 (Wang et al., 2022) 571.1 / 67.3 56.3 / 66.2
@@
2-shot S3HQA GPT 3.5 direct (Lei et al., 2023) 57.1 / 68.8 -
2-shot S3HQA GPT 3.5 CoT (Lei et al., 2023) 60.3 / 72.1 -
2-shot Baseline Standard prompting LLaMA 3-70B 48.97 / 60.18 52.88 / 61.42
2-shot Baseline CoT LLaMA 3-70B 54.22 / 64.98 55.71 / 62.24
2-shot TTQA-RS LLaMA 3 - 70B 63.12 / 73.61 61.97 / 67.56
2-shot TTQA-RS LLaMA 3 - 70B + GPT-4 65.49 / 76.43 63.69 / 71.83
@@
(a) EM score
(b) F1 score
Figure 4: Performance of HybridQA test set on different LLaMA models
(a) EM score
(b) F1 score
Figure 5: Performance of OTT-QA dev set on different LLaMA models
@@
2-shot Baseline Standard prompting LLaMA3-70B 53.28 59.65
2-shot Baseline CoT LLaMA3-70B 55.74 63.50
2-shot TTQA-RS LLaMA 3 - 70B 63.15 70.84
2-shot TTQA-RS LLaMA 3 - 70B + GPT-4 67.27 79.55
@@
Baseline Standard prompting 48.97 / 60.18 52.88 / 61.24 53.28 / 59.65
Baseline CoT 54.22 / 64.98 55.71 / 62.24 55.74/63.50
CoT + entity-type prediction of expected answer 59.68 / 67.47 57.75 / 64.65 57.53 / 66.24
@@
Question decomposition and including sub-question with generated sub-answer + entity-type prediction of expected answer with COT (no summarization) 61.14 / 70.45 58.72 / 66.40 61.23 / 69.14
@@
Our model - Question decomposition and including sub-question with generated sub-answer + entity-type prediction of expected answer + summarization with CoT 63.12 / 73.61 61.97 / 67.56 63.15 / 70.84
Table 4: Human evaluation of generated summaries for a sampled test set of HybridQA
Human Evaluation Metrics Performance
Correctness 0.94
Inclusivity 0.98
Completeness 0.71
@@
In this section, we discuss all our major findings. Table 1 displays the performance of our model with other existing models on the HybridQA dataset. We used exact match (EM) and F1 score to evaluate the performance the table-text QA models. From Table 1, we can observe that most existing models train their models for table-text QA. S3HQA (Lei et al., 2023) is the only model among the existing works that uses in-context learning for the HybridQA dataset. Please note that "S3HQA GPT-3.5 direct" refers to S3HQA model with standard prompting using GPT-3.5. Our TTQA-RS model with LLaMA 3-70B on the HybridQA’s development set was able to beat S3HQA’s CoT with GPT 3.5 by 3% exact match. Our 2-shot model with LLaMA-4 also beats the baseline standard and CoT prompting models by a huge margin (i.e. by 9% exact match when compared with standard prompting and by 6% exact match when compared with CoT in the test set). Furthermore, in Figure 4 and in Figure 5, we have shown the performance of our TTQA-RS model on different parameter models of LLaMA 2 and LLaMA 3 models on HybridQA test set and the OTT-QA development set respectively. For all the different parameter models of LLaMA-2 and LLaMA-3, our framework performed better than the baseline prompting models (i.e. standard prompting and CoT prompting). Our experiments show that our breakdown prompting approach with summarization and reasoning can improve the performance of all open-source models for table-text QA tasks. To show that our TTQA-RS approach can improve table-text QA on also GPT model, we have experimented with GPT-4 in the last stage of our model, i.e. table-text QA on the original question. For the remaining stages of the reader, we have used LLaMA 3- 70b. By adding GPT-4 in the last step, we were able to show the best performance with an exact match of 65.49 and F1 score of 76.43 in the development set, and an exact match of 63.69 and F1 score of 71.83 on the test set. With 2-shot learning using TTQA-RS LLaMA 3 70B + GPT-4 model, we were able to reach a model performance very close to the best existing training-based model (S3HQA with supervised learning) on the HybridQA dataset. For cost limitations, we have limited experiments with GPT-4 to only the last stage of our model.
@@
Table 2 shows the performance of our model - TTQA-RS on the OTT-QA development set. To the best of our knowledge, MFORT-QA (Guan et al., 2024) is the only model that has used in-context learning for the OTT-QA dataset, but since they have not reported their performance on the development set, we therefore compare our model’s performance with other existing works that trained the models. Our TTQA-RS model with LLaMA 3-70B + GPT-4 model achieved the best performance (exact match of 67.27 and F1 score of 79.55) on the development set and has achieved new state-of-the-art performance of the OTT-QA’s development set.
With the evaluation of our model - TTQA-RS on the HybridQA and OTT-QA development set, we have shown the potential of prompting approaches with small language models (like LLaMA) and also using GPT-4.

### v2

@@ Abstract
Question answering (QA) over tables and text has gained much popularity over the years. Multi-hop table-text QA requires multiple hops between the table and text, making it a challenging QA task. Although several works have attempted to solve the table-text QA task, most involve training the models and requiring labeled data. In this paper, we have proposed a Retrieval Augmented Generation (RAG) based model - “TTQA-RS: A break-down prompting approach for Multi-hop Table-Text Question Answering with Reasoning and Summarization” [^1] Link to code in Github: https://github.com/jayetri/TTQA-RS. Our model uses an enhanced retriever for table-text information retrieval and uses augmented knowledge, including table-text summary with decomposed sub-questions with answers for a reasoning-based table-text QA. Using open-source language models, our model outperformed all existing prompting methods for table-text QA tasks on existing table-text QA datasets, such as HybridQA and OTT-QA’s development set. Our experiments demonstrate the potential of prompt-based approaches using open-source LLMs. Additionally, by using LLaMA3-70B, our model achieved state-of-the-art performance for prompting-based methods on multi-hop table-text QA.
@@ 1 Introduction
Over the years, several works have attempted to solve this task. But the majority of these works have used supervised-training for the reader, requiring a large amount of labeled data (Chen et al., 2020b; Sun et al., 2021; Wang et al., 2022b; Eisenschlos et al., 2021; Feng et al., 2022; Kumar et al., 2023; Chen et al., 2020a; Li et al., 2021). In this paper, we have proposed a prompting-based approach while using open-source large language models (LLMs) for multi-hop table-text QA.
@@
In this paper, we introduced a framework - TTQA-RS, a prompting approach for table-text QA that despite CoT’s shortcomings on small-parameter models, we were able to reduce the hallucinations on open-source small models. Furthermore, our proposed model was able to beat the state-of-the-art model - S3HQA’s CoT prompting using different LLaMA models (Lei et al., 2023) on HybridQA dataset. By beating their model’s performance, we have shown the potential for smaller LLMs in multi-hop table-text QA.
For our experiments, we have used HybridQA dataset and OTT-QA’s development set. OTT-QA is an extension of the HybridQA dataset. Similar to the HybridQA dataset, the OTT-QA dataset is also constructed using questions based on Wikipedia tables and text. But unlike the HybridQA dataset, the test set of the OTT-QA dataset does not have hyperlinks in the table cells that can be linked to the Wikipedia text. Hence, the OTT-QA’s test set is more challenging. The task of linking tables and text passages for open-domain QA is out of scope of this paper.
The TTQA-RS model breaks down the table-text QA problem into multiple steps. In the HybridQA and the OTT-QA dataset, the questions require multiple steps of reasoning over table and text to answer. The TTQA-RS model generates the sub-questions that can help in answering the complex questions. It also generates the summary of the table and text, which is in turn used for the table-text QA of the original questions. Breaking down the complex multi-hop QA problem into simple, smaller steps along with providing augmented information, including the table-text summary, can improve the performance of multi-hop table-text QA tasks using small open-source LLMs. We have also proposed an enhanced retriever for our framework that further improves the overall performance of our table-text QA model. In Figure 1, we show an example of a question that uses standard prompting, CoT, and the TTQA-RS approach for multi-hop QA.
@@ 2 Related Works
Multi-hop table-text QA can be a complex task as it requires multiple hops between the table and text to answer the questions. S3HQA (Lei et al., 2023) and MFORT-QA (Guan et al., 2024) are the only two existing models as per our knowledge that use in-context learning for multi-hop table-text QA. The S3HQA model has demonstrated table-text QA task using the Hybrid-QA dataset, whereas MFORT-QA has used the OTT-QA dataset. The S3HQA model uses a three-step method - a retriever with refinement training, a hybrid selector, and a generation-based reasoner with GPT 3.5 for the hybrid table-text QA task. MFORT-QA uses the Chain-of-thought (CoT) method to break down complex questions into smaller sub-questions, and uses Retrieval Augmented Generation to extract more context. Similar to the MFORT-QA model, we also use a RAG-based framework and break down complex questions into smaller sub-questions. With the complexity of the multi-hop QA task broken down into smaller questions, LLMs are in turn working on a smaller problem and perform better as single-step reasoners. Our model - TTQA-RS, additionally uses an enhanced retriever for table-text information retrieval and further generates a summary using the retrieved table rows and passages. Then, for table-text question answering (QA), it uses the generated summary, the predicted entity type of the answer, and the generated sub-questions along with the answer.
@@ 3.1 System Overview
Figure 2: An overview of TTQA-RS framework. The brown dashed lines represent the retriever and the black dotted lines represent the reader for the table-text QA model.
The TTQA-RS model uses a RAG-based retriever-reader framework. Our reader breaks down the table-text QA problem into five steps - (1) Summary generation using retrieved tables rows and passages, (2) Question decomposition, (3) Entity type prediction of the expected answer, (4) Table-text QA of independent sub-question, and (5) Table-text QA of the original question. Figure 2 shows an overview of the TTRS-QA framework.
@@ 3.2 Retriever
Figure 3: Example of rows to sentences for a table
We have used an enhanced retriever that first converts rows in the tables to sentences using a pre-trained BART-Large model. The questions and sentences are encoded and then fed to a fine-tuned Dense Passage Retriever (DPR) (we can call this the table retriever). The table retriever retrievers the top 3 most relevant rows. It is important to first convert the rows to sentences using BART since the retriever is not always smart enough to interpret the rows for retrieval. This step helps the retriever to make more intelligent decisions. Figure 3 shows an example where the rows of the table are converted to sentences. In this example, when the table is directly fed to the retriever, it fails to understand that rank 2 implies that the player came second in the National Football League. When the rows are converted to sentences before feeding to the DPR, it can better understand the table information and, in turn, retrieve more accurate information.
For text retrieval, we extract the passages from the text linked to the top 3 rows using hyperlinks. The question embedding and row embedding refer to the final dense vectors generated from the encoding process of the question and the rows, respectively. We calculate the combined embeddings using the following:
[…]
Figure 4: Example of our approach using TTQA-RS model
In the Equation 3.2, $\alpha$ is a parameter with values between 0 and 1. We estimate the similarity between the combined embeddings and the text embeddings to retrieve the top 6 candidates using a pre-trained DPR model (i.e. the text retriever). These top 6 candidates, along with the question keywords generated using a pre-trained T5 model, are fed into a re-ranker, (a pre-trained ’deepset/roberta-base-squad2’ model). The re-ranker identifies the top 3 candidates, which are subsequently forwarded to the reader.
@@ 3.3.1 Table-text Summarization
This is the first step of the reader model. The retrieved rows and passages are used to generate summaries of the table and text. We used zero-shot learning with LLaMA 3-70B model to generate the summaries. In Appendix B we have shown an example of a table-text summarization prompt.
@@ 3.3.2 Question decomposition
In the next step, we break down the questions and identify the sub-questions, such that the answer of one sub-question can aid in answering the original complex question. From here onwards, we will refer to the sub-question that can be answered first as the “independent sub-question”. Let’s take the example in Figure 4. The complex question - "What was the release date of the game which Andrew Voss provided commentary on ?" can be broken down into sub-questions. The independent sub-question for this question is - "Which game has Andrew Voss provided commentary on?". The answer to this sub-question is "Rugby League 3". This can be used to simplify the original complex question to the following - " What was the release date of Rugby League 3?". Thus, including the information about the independent sub-question and the sub-answer helps to reduce the complexity of the multi-hop task. Identifying the independent sub-question and breaking down the complex multi-hop QA problem helps to reduce the complexity of the problem, and in turn, boosts the accuracy of the model. We use in-context learning with LLaMA3-70B model to generate the independent sub-questions for the given complex queries. Appendix C shows further examples.
@@ 4.1 Datasets
OTT-QA (Chen et al., 2020a) is an open-domain multi-hop table-text QA dataset. For our experiments, we only use the development set of the OTT-QA dataset.
@@ 4.3 Baseline Models
Standard prompting - For the baseline standard prompting model, we used the same retriever as in TTQA-RS model. For the reader, we performed in-context learning with standard prompting (Brown et al., 2020) for the QA task.
Chain of Thought Prompting (CoT) - The CoT baseline model also uses the same retriever as the TTQA-RS model. The reader uses in-context learning with CoT prompting Wei et al. (2022).
Least-to-Most Prompting - Least-to-Most Prompting (LtM) Zhou et al. (2022) extends CoT prompting by first diving a problem into subproblems and then solving each one. For the baseline LtM model, we use TTQA-RS’s retriever with LtM prompting Zhou et al. (2022).
CoT with Self Consistency (sc) - Wang et al. (2022a) proposed self-consistency as a replacement for the naive greedy decoding utilized in CoT prompting.
@@ 5.1 Main Results
Table 1: Retrieval Accuracy (%)
Metrics HybridQA OTT-QA
Dev Test Dev
HIT@1 73.28 71.49 72.53
HIT@3 82.01 80.65 81.83
@@
Train MuGER2 (Wang et al., 2022b) 57.11 / 67.3 56.3 / 66.2
@@
2-shot S3HQA LLaMA 3 - 70B CoT (Lei et al., 2023) 52.23 / 58.49 53.75 / 62.94
2-shot Baseline Standard prompting LLaMA 3-70B 48.73 / 60.02 50.76 / 59.23
2-shot Baseline CoT LLaMA 3-70B 54.53 / 62.49 55.89 / 62.98
2-shot Baseline CoT LLaMA 3-70B (with sc) 54.63 / 64.07 55.96 / 63.07
2-shot Baseline Least-to-Most prompting LLaMA 3-70B 57.84 / 65.42 57.83 / 67.39
2-shot TTQA-RS LLaMA 3 - 70B 63.97 / 74.20 62.34 / 70.58
2-shot TTQA-RS LLaMA 3 - 70B (with sc) 63.98 / 74.32 62.92 / 70.83
@@
Figure 5: EM-score of HybridQA test set on different LLaMA models
@@
2-shot Baseline Standard prompting LLaMA3-70B 51.24 58.73
2-shot Baseline CoT LLaMA3-70B 55.78 64.30
2-shot Baseline CoT LLaMA3-70B (with sc) 55.79 65.54
2-shot Baseline Least-to-Most prompting LLaMA3-70B 58.68 67.75
2-shot TTQA-RS LLaMA 3 - 70B 63.17 73.52
2-shot TTQA-RS LLaMA 3 - 70B (with sc) 63.31 73.54
@@
Baseline Standard prompting 48.73/60.02 50.76/59.23 51.24/58.73
Baseline CoT 54.53/62.49 55.89/62.98 55.78/64.30
CoT + entity-type prediction of expected answer 55.34/63.95 56.52/63.47 57.60/67.84
@@
Question decomposition and including sub-question with generated sub-answer + entity-type prediction of expected answer with COT (no summarization) 60.85/70.68 58.32/66.82 61.06/70.48
@@
Our model - Question decomposition and including sub-question with generated sub-answer + entity-type prediction of expected answer + summarization with CoT 63.97/74.20 62.34/70.58 63.17/73.52
@@
In this section, we discuss all our major findings. Table 1 shows the hit score of our model on the HybridQA and OTT-QA dataset. Table 2 displays the performance of our model with other existing models on the HybridQA dataset. From Table 2, we can observe that most existing models train their models for table-text QA. S3HQA (Lei et al., 2023) is the only model among the existing works that uses in-context learning for the HybridQA dataset. Our TTQA-RS model with LLaMA 3-70B on the HybridQA’s development set was able to beat S3HQA’s CoT model by 11.7% exact match (EM) and by 15.71% F1 score. Our model outperformed the S3HQA-CoT model on the HybridQA test using LLaMA3-70B by 8.6% and 7.6% exact match (EM) and F1 score respectively. Our 2-shot model with LLaMA 3-70B also beats the baseline standard, baseline CoT, baseline COT with sc, and Least-to-Most prompting method by a huge margin. This shows the effectiveness of our reader model. We can also observe that our baseline CoT model outperformed the S3HQA’s CoT model for the HybridQA dataset. This shows the effectiveness of our retriever model. Furthermore, in Figure 5, we have shown the performance of our TTQA-RS model on different parameter models of LLaMA 2 and LLaMA 3 models on HybridQA test set. For all the different parameter models of LLaMA-2 and LLaMA-3, our framework performed better than the baseline prompting models and also the S3HQA-CoT model. Our experiments show that our approach can improve the performance of open-source models for table-text QA tasks.
@@
Table 3 shows the performance of our model - TTQA-RS on the OTT-QA development set. To the best of our knowledge, MFORT-QA (Guan et al., 2024) is the only model that has used in-context learning for the OTT-QA dataset, but since they have not reported their performance on the development set, we therefore compare our model’s performance with other existing works that trained the models. Our TTQA-RS model with LLaMA 3-70B with self-consistency has achieved new state-of-the-art performance on the OTT-QA’s development set.